# PersonalView: Multi-View Consistent Human Image Customization via In-Context Learning

### v1

@@ 4 Experiments
4.1 Experimental Setup
Implementation details. We utilize the recently developed DiT model FLUX (Labs, 2024b) with a pre-trained personalized module PuLID (Guo et al., 2024) as our foundational model. Our training set comprises only 100 cases randomly sampled from NeRSemble (Kirschstein et al., 2023), a widely utilized multi-view human dataset. Please refer to the Appendix for more details.
Baselines. To the best of our knowledge, our approach is the first to support end-to-end customized multi-view human image generation. For comparison, we construct a two-stage baseline wherein PuLID is used for identity-specific image synthesis, followed by the novel view synthesis method based on image-to-3D reconstruction. Specifically, we benchmark the proposed method against DiffPortrait3D (Gu et al., 2024) and Era3D (Li et al., 2024d), recent methods synthesizing 3D-consistent photo-realistic novel views. We also compare with ViewCrafter (Yu et al., 2024), a recent method synthesizing high-fidelity novel views of generic scenes. Besides, we compare with recent unified multi-modal model BAGEL (Deng et al., 2025) and Qwen-Image (Wu et al., 2025).
Evaluation. We collect a diverse portrait test set from the internet which consists of 80 characters ensuring demographic representation, with 50 prompts for comprehensive motion evaluation. To assess multi-view consistency, we follow the 3D consistency metric (MV Cons.) proposed in Pippo (Kant et al., 2025). MV Cons. is calculated on a pair of images where landmarks are estimated first and then reprojected to each other for error calculation. More details are in Appendix. The evaluation framework employs CLIP-T and CLIP-I for quantitative assessment of text and image alignment. To further measure identity preservation, we implement facial recognition embedding similarity metrics (An et al., 2021) complemented by specialized facial motion analysis protocols.
4.2 Main Results
Qualitative Results. We present a qualitative assessment comparing PersonalView with baseline methods. As illustrated in Figure 4, DiffPortrait3D struggles with maintaining full-body geometric consistency and coherent scene appearance, largely due to its design and training focus on portrait-level generation. Similarly, Era3D, by primarily targeting subject-centric multi-view synthesis, performs poorly in preserving visual consistency across scene elements. Although ViewCrafter achieves improved scene coherence, it does so at the cost of human body geometry, often leading to distorted multi-view results. For BAGEL and Qwen-Image, they both exhibit relatively poor performance in multi-view control and consistency. In contrast, our proposed method achieves strong geometric consistency for the entire human body, along with visually consistent scene modeling, including accessories and backgrounds, while simultaneously preserving identity fidelity and supporting prompt-guided customization. More results are included in Figure 6 of the Appendix.
@@ 4.2 Main Results
PersonalView 31.80 26.92 29.67 24.56
Table 2: User Study. Our PersonalView achieves the best human preference compared with all baselines.
Quantitative Results. We report the results of the quantitative comparison in Table 3. As observed, DiffPortrait3D exhibits limited multi-view consistency, highlighting its weakness in modeling coherent geometry across viewpoints. Although Era3D mitigates this issue to some extent, it still suffers from subpar performance in identity preservation and semantic control accuracy. Likewise, ViewCrafter shows inferior results in both identity consistency and CLIP-based semantic alignment. Although BAGEL and Qwen-Image improve multi-view consistency, the identity similarity has decreased. In contrast, our method consistently achieves superior performance across all key evaluation dimensions, including multi-view consistency, identity fidelity, and prompt-aligned semantic controllability.
DiffPortrait3D (Gu et al., 2024) 7.887 0.7721 0.2588 0.6689
Era3D (Li et al., 2024d) 6.383 0.6462 0.2421 0.6708
ViewCrafter (Yu et al., 2024) 6.945 0.5820 0.2358 0.7037
BAGEL (Deng et al., 2025) 5.882 0.5623 0.2591 0.7289
Qwen-image (Wu et al., 2025) 5.324 0.6433 0.2603 0.7422
PersonalView 3.697 0.7920 0.2615 0.7793
@@
User Study. To further evaluate the effectiveness of our methodology, we conduct a human-centric assessment, comparing our approach with existing novel view synthesis methods. We recruit 25 evaluators to assess 40 sets of generated results. For each set, we present reference images alongside multi-view images produced using identical seeds across various methods. The quality of the generated multi-view images is evaluated based on four criteria: Multi-View Consistency, Text Alignment, ID Consistency, and Overall Quality. As depicted in Table 2, our PersonalView achieves higher user preference across all evaluative dimensions, underscoring its superior effectiveness.
@@ 4.3 Ablation Study
Multi-View In-Context Learning. To validate the effectiveness of our proposed components, we conduct an ablation study presented in Figure 5 and Table 4. As shown in the figure, the pretrained model often generates head orientations inconsistent with the corresponding depth maps, leading to degraded geometric consistency across viewpoints. Incorporating multi-view in-context learning significantly alleviates this issue, resulting in more coherent human geometry across views. However, due to the limited diversity of scenes in the multi-view dataset, prompt-based semantic controllability, particularly for attributes such as background and clothing, tends to diminish. In contrast, Semantic Correspondence Alignment (SCA) loss effectively preserves fine- grained semantic control, ensuring that identity and prompt-relevant attributes are maintained throughout the generated views. More ablation studies are given in the Appendix.
Number of Training Iterations and Samples. To investigate the impact of training iterations on model performance, we present a quantitative comparison in Table 5(a). As indicated by the results, 800 iterations provide a favorable trade-off between efficiency and performance, and are therefore adopted in our final setting. Additionally, to validate the lightweight nature of our in-context learning framework, we evaluate the effect of varying the number of training samples. As shown in Table 5(b), our method achieves competitive performance with as few as 100 training samples, demonstrating its data efficiency and strong generalization capability under limited supervision.
5 Conclusion
In this work, we introduce PersonalView, a novel framework for multi-view human image customization from a single photograph. By leveraging in-context learning with multi-view depth maps, our method enhances DiT-based models’ multi-view reasoning. We also propose a Semantic Correspondence Alignment Loss to preserve prompt controllability during fine-tuning. Extensive experiments show that PersonalView excels in identity fidelity and coherent multi-view synthesis, requiring minimal additional training. Our approach enables efficient, high-quality multi-view customization without large-scale multi-view datasets.
Appendix A Appendix
A.1 Implementation Details
For the NeRSemble dataset, we employ Qwen-VL (Bai et al., 2023) to generate descriptive captions. During training, we learn the LoRA for 800 iterations with a learning rate of 1e-4 with batch size 1. We employ the AdamW optimizer with a weight decay parameter of 1e-2. The epsilon is set to the default 1e-8 and the weight decay is set to 1e-2. During inference, we use 50 steps of DDIM sampler and classifier-free guidance with a scale of 7.5. We generate multi-view images with 512 × 512 spatial resolution. We used Lora rank 128, following the common setting of FLUX finetuning. All experiments are conducted on a single NVIDIA A800 GPU. Our code will be open-source.
A.2 Evaluation Metric for Multi-View Consistency
To evaluate the multi-view consistency of the generated results, we adopt the re-projection error metric proposed in Pippo (Kant et al., 2025). Specifically, we first estimate facial landmarks from the generated images and then establish pairwise correspondences of these landmarks across different views. Based on these correspondences, we apply Triangulation using the Direct Linear Transformation algorithm to recover the 3D positions of each landmark. Finally, we reproject the 3D landmarks back onto each view and compute the Reprojection Error as the L2 distance between the original 2D landmark and the reprojected point, normalized by the image resolution. The final RE score is obtained by averaging this error across all views and landmarks.
Method MV Cons. ($\downarrow$) CLIP-T ($\uparrow$) ID Cons. ($\uparrow$)
ImagPose 4.882 0.2385 0.7315
MVDream 8.669 0.2322 0.7286
Ours 3.697 0.2604 0.7920
@@ A.2 Evaluation Metric for Multi-View Consistency
A.3 More Ablation Study
The core design of SCA is to preserve the original model’s semantic control (text-image correspondence) without compromising in-context learning during fine-tuning. In contrast, applying regularization on other aspects like full-attention features or model parameters leads to reduced multi-view consistency as shown in Table 7, since it disrupts in-context learning.
@@ A.3 More Ablation Study
Table 7: Quantitative ablation study of different alignment loss.
A.4 Comparison with more baselines
We provide comparisons with additional potential baselines, such as ImagPose (Shen & Tang, 2024) and MVDream (Shi et al., 2023b). Compared with ImagPose, our method enables the customization model to generate multi-view consistent images via lightweight adaptation with only 100 samples, while Imagpose requires  85000 images with diverse viewpoints. Besides, Imagpose lacks the ability to control rich semantics, such as background changes via prompts, while our method preserves the precise semantic control as validated in Table 6. For MVDream, it exhibits limited multiview consistency and semantic control, probably because it has difficulty generalizing to the human domain.
A.5 Discussion of SMPL Fitting
@@
Figure 6: More results of PersonalView.
A.6 Difference from IC-LoRA.
In practice, IC-LoRA (Huang et al., 2024) boils down to fine-tuning the DiT on concatenated images, which is a simple yet widely-applicable technique. However, to our knowledge, there is no multi-view conditional generation model built on IC-LoRA at the time of submission. Our work provides solid evidence of the success of in-context learning in transferring multi-view generation to a generic model. On the other hand, our experiments demonstrate the insufficiency of using in-context learning alone, which has distinct contributions not covered by IC-LoRA.
A.7 Limitation
While our approach demonstrates strong performance in generating multi-view consistent customized images, it also has certain limitations. In particular, it primarily relies on activating the in-context learning capabilities of the base model, and therefore, its effectiveness is inherently constrained by the representational power of the underlying pre-trained model. Nevertheless, a key advantage of our framework lies in its modularity and transferability, making it easily adaptable to more powerful backbone models as they emerge — a direction we plan to explore in future work.
A.8 Ethics Statement
@@
A.9 Reproducibility Statement
We make the following efforts to ensure the reproducibility of PersonalView: (1) Our training and inference codes together with the trained model weights will be publicly available. (2) We provide training details in the appendix, which is easy to follow. (3) We provide the details of the human evaluation setups.
A.10 LLM Usage Statement

### v2

@@ 4 Experiments
DiffPortrait3D [gu2024diffportrait3d] 7.887 0.7721 0.2588 0.6689
Era3D [li2024era3d] 6.383 0.6462 0.2421 0.6708
ViewCrafter [yu2024viewcrafter] 6.945 0.5820 0.2358 0.7037
BAGEL [deng2025emerging] 5.882 0.5623 0.2591 0.7289
Qwen-image [wu2025qwen] 5.324 0.6433 0.2603 0.7422
MagicView 3.697 0.7920 0.2615 0.7793
Table 3: User Study. Our MagicView achieves the best human preference compared with all baselines.
@@
MagicView 31.80 26.92 29.67 24.56
4.1 Experimental Setup
Implementation details. We utilize the recently developed DiT model FLUX [flux1-dev] with a pre-trained personalized module PuLID [guo2024pulid] as our foundational model. Our training set comprises only 100 cases randomly sampled from NeRSemble [kirschstein2023nersemble], a widely utilized multi-view human dataset. Please refer to the Appendix for more details.
Baselines. To the best of our knowledge, our approach is the first to support end-to-end customized multi-view human image generation. For comparison, we construct a two-stage baseline wherein PuLID is used for identity-specific image synthesis, followed by the novel view synthesis method based on image-to-3D reconstruction. Specifically, we benchmark the proposed method against DiffPortrait3D [gu2024diffportrait3d] and Era3D [li2024era3d], recent methods synthesizing 3D-consistent photo-realistic novel views. We also compare with ViewCrafter [yu2024viewcrafter], a recent method synthesizing high-fidelity novel views of generic scenes. Besides, we compare with recent unified multi-modal model BAGEL [deng2025emerging] and Qwen-Image [wu2025qwen].
@@ 4.1 Experimental Setup
Evaluation. We collect a diverse portrait test set from the internet which consists of 80 characters ensuring demographic representation, with 50 prompts for comprehensive motion evaluation. To assess multi-view consistency, we follow the 3D consistency metric (MV Cons.) proposed in Pippo [kant2025pippo]. MV Cons. is calculated on a pair of images where landmarks are estimated first and then reprojected to each other for error calculation. More details are in Appendix. The evaluation framework employs CLIP-T and CLIP-I for quantitative assessment of text and image alignment. To further measure identity preservation, we implement facial recognition embedding similarity metrics [an2021partial] complemented by specialized facial motion analysis protocols.
4.2 Main Results
Qualitative Results. We present a qualitative assessment comparing MagicView with baseline methods. As illustrated in Figure 4, DiffPortrait3D struggles with maintaining full-body geometric consistency and coherent scene appearance, largely due to its design and training focus on portrait-level generation. Similarly, Era3D, by primarily targeting subject-centric multi-view synthesis, performs poorly in preserving visual consistency across scene elements. Although ViewCrafter achieves improved scene coherence, it does so at the cost of human body geometry, often leading to distorted multi-view results. For BAGEL and Qwen-Image, they both exhibit relatively poor performance in multi-view control and consistency. In contrast, our proposed method achieves strong geometric consistency for the entire human body, along with visually consistent scene modeling, including accessories and backgrounds, while simultaneously preserving identity fidelity and supporting prompt-guided customization. More results are included in the Appendix.
Quantitative Results. We report the results of the quantitative comparison in Tab. 2. As observed, DiffPortrait3D exhibits limited multi-view consistency, highlighting its weakness in modeling coherent geometry across viewpoints. Although Era3D mitigates this issue to some extent, it still suffers from subpar performance in identity preservation and semantic control accuracy. Likewise, ViewCrafter shows inferior results in both identity consistency and CLIP-based semantic alignment. Although BAGEL and Qwen-Image improve multi-view consistency, the identity similarity has decreased. In contrast, our method consistently achieves superior performance across all key evaluation dimensions, including multi-view consistency, identity fidelity, and prompt-aligned semantic controllability.
Table 4: Quantitative comparison with ImagPose and MVDream.
Method MV Cons. ($\downarrow$) CLIP-T ($\uparrow$) ID Cons. ($\uparrow$)
ImagPose 4.882 0.2385 0.7315
MVDream 8.669 0.2322 0.7286
Ours 3.697 0.2604 0.7920
@@ 4.2 Main Results
4.3 Comparison with More Methods
We provide comparisons with additional potential baselines, such as ImagPose [shen2024imagpose] and MVDream [shi2023mvdream]. Compared with ImagPose, our method enables the customization model to generate multi-view consistent images via lightweight adaptation with only 100 samples, while Imagpose requires  85000 images with diverse viewpoints. Besides, Imagpose lacks the ability to control rich semantics, such as background changes via prompts, while our method preserves the precise semantic control as validated in Tab. 4. For MVDream, it exhibits limited multiview consistency and semantic control, probably because it has difficulty generalizing to the human domain.
User Study. To further evaluate the effectiveness of our methodology, we conduct a human-centric assessment, comparing our approach with existing novel view synthesis methods. We recruit 25 evaluators to assess 40 sets of generated results. For each set, we present reference images alongside multi-view images produced using identical seeds across various methods. The quality of the generated multi-view images is evaluated based on four criteria: Multi-View Consistency, Text Alignment, ID Consistency, and Overall Quality. As depicted in Tab. 3, our MagicView achieves higher user preference across all evaluative dimensions, underscoring its superior effectiveness.
@@ 4.4 Ablation Study
Multi-View In-Context Learning. To validate the effectiveness of our proposed components, we conduct an ablation study presented in Figure 5 and Tab. 5. As shown in the figure, the pretrained model often generates head orientations inconsistent with the corresponding depth maps, leading to degraded geometric consistency across viewpoints. Incorporating multi-view in-context learning significantly alleviates this issue, resulting in more coherent human geometry across views. However, due to the limited diversity of scenes in the multi-view dataset, prompt-based semantic controllability, particularly for attributes such as background and clothing, tends to diminish. In contrast, Semantic Correspondence Alignment (SCA) loss effectively preserves fine- grained semantic control, ensuring that identity and prompt-relevant attributes are maintained throughout the generated views. More ablation studies are given in the Appendix.
Number of Training Iterations and Samples. To investigate the impact of training iterations on model performance, we present a quantitative comparison in Tab. 6(a). As indicated by the results, 800 iterations provide a favorable trade-off between efficiency and performance, and are therefore adopted in our final setting. Additionally, to validate the lightweight nature of our in-context learning framework, we evaluate the effect of varying the number of training samples. As shown in Tab. 6(b), our method achieves competitive performance with as few as 100 training samples, demonstrating its data efficiency and strong generalization capability under limited supervision.
Design of SCA. The core design of SCA is to preserve the original model’s semantic control (text-image correspondence) without compromising in-context learning during fine-tuning. In contrast, applying regularization on other aspects like full-attention features or model parameters leads to reduced multi-view consistency as shown in Tab. 7, since it disrupts in-context learning.
@@
5 Conclusion
In this work, we introduce MagicView, a novel framework for multi-view human image customization from a single photograph. By investigating and leveraging in-context learning with multi-view depth priors, our method enhances DiT-based models’ multi-view reasoning. We also propose a Semantic Correspondence Alignment Loss to preserve prompt controllability during fine-tuning. Extensive experiments show that MagicView excels in identity fidelity and coherent multi-view synthesis, requiring minimal additional training. Our approach enables efficient, high-quality multi-view customization without large-scale multi-view datasets.
6 Limitation
While our approach demonstrates strong performance in generating multi-view consistent customized images, it also has certain limitations. In particular, it primarily relies on activating the in-context learning capabilities of the base model with multi-view priors, and therefore, its effectiveness is inherently constrained by the representational power of the underlying pre-trained model. Nevertheless, a key advantage of our framework lies in its modularity and transferability, making it easily adaptable to more powerful backbone models as they emerge, a direction we plan to explore in future work.
[…]
Supplementary Material
7 Implementation Details
For the NeRSemble dataset, we employ Qwen-VL [bai2023qwenvlversatilevisionlanguagemodel] to generate descriptive captions. During training, we learn the LoRA for 800 iterations with a learning rate of 1e-4 with batch size 1. We employ the AdamW optimizer with a weight decay parameter of 1e-2. The epsilon is set to the default 1e-8 and the weight decay is set to 1e-2. During inference, we use 50 steps of DDIM sampler and classifier-free guidance with a scale of 7.5. We generate multi-view images with 512 × 512 spatial resolution. We used Lora rank 128, following the common setting of FLUX finetuning. All experiments are conducted on a single NVIDIA A800 GPU. Our code will be open-source.
8 Evaluation Metric for Multi-View Consistency
To evaluate the multi-view consistency of the generated results, we adopt the re-projection error metric proposed in Pippo [kant2025pippo]. Specifically, we first estimate facial landmarks from the generated images and then establish pairwise correspondences of these landmarks across different views. Based on these correspondences, we apply Triangulation using the Direct Linear Transformation algorithm to recover the 3D positions of each landmark. Finally, we reproject the 3D landmarks back onto each view and compute the Reprojection Error as the L2 distance between the original 2D landmark and the reprojected point, normalized by the image resolution. The final RE score is obtained by averaging this error across all views and landmarks.
9 Discussion of SMPL Fitting
@@
Figure 6: More results of MagicView.
10 Difference from IC-LoRA.
In practice, IC-LoRA [lhhuang2024iclora] boils down to fine-tuning the DiT on concatenated images, which is a simple yet widely-applicable technique. However, to our knowledge, there is no multi-view conditional generation model built on IC-LoRA at the time of submission. Our work provides solid evidence of the success of in-context learning in transferring multi-view generation to a generic model. On the other hand, our experiments demonstrate the insufficiency of using in-context learning alone, which has distinct contributions not covered by IC-LoRA.
11 Ethics Statement
@@
12 Reproducibility Statement
We make the following efforts to ensure the reproducibility of MagicView: (1) Our training and inference codes together with the trained model weights will be publicly available. (2) We provide training details in the appendix, which is easy to follow. (3) We provide the details of the human evaluation setups.
13 LLM Usage Statement